OpenStudy (anonymous):

MEDALS N BECOMING FAN 1.)Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1 -6, -12, -24, -48, -96, -192 -3, -6, -12, -24, -48, -96 0, 2, -6, -4, -2, 0 -3, -6, -4, -2, 0, 2 2.)Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ... an = 2 ● 4^(n+1) an = 2 ● (-4)^n an = 2 ● 4^(n-1 ) an = 2 ● (-4)^(n-1)

OpenStudy (danjs):

The first sequence, is taking the term before it and multiplying it by 2 every time

OpenStudy (danjs):

starting with -3

OpenStudy (danjs):

\[a _{2}=2 * a_{2-1}\] a2 would be 2 times a1 a3 would be 2 times a2 a4 would be 2 times a3 . . .

OpenStudy (anonymous):

-6, -12, -24, -48, -96, -192

OpenStudy (solomonzelman):

but your first term is?

OpenStudy (anonymous):

#2 i got an = 2 ● 4^(n-1 )?

OpenStudy (anonymous):

-3, -6, -12, -24, -48, -96

OpenStudy (anonymous):

sry i meant that 1

OpenStudy (solomonzelman):

yeah that is better

OpenStudy (solomonzelman):

and for 2 u got?

OpenStudy (anonymous):

is my answer for 2 correct?

OpenStudy (solomonzelman):

you are multiplying times what every time?

OpenStudy (danjs):

Since they gave you multiple choice, you can try each one and really know nothing..

OpenStudy (solomonzelman):

(the form is a\(_{\rm n}\)=(a\(_{\rm 1}\))•r\(^{\rm n-1}\) ) r - common ratio - (the number by which you multiply each time)

OpenStudy (solomonzelman):

in this case: 2, -8, 32, -128 you are multiplying times 4? That would be close, but not exactly right, because your values are changing signs (they alternate)

OpenStudy (solomonzelman):

2 • what = -8? -8• what = 32

OpenStudy (anonymous):

an = 2 ● (-4)^n @SolomonZelman

OpenStudy (solomonzelman):

ok, i am not going to tell you right or wrong, but this is what I will say \(a_1=2\cdot (-4)^{1}\) \(a_1=2\cdot(-4)\) \(a_1=-8\)

OpenStudy (solomonzelman):

but -8 is the second term, not the first

OpenStudy (anonymous):

oh so it has to be this one an = 2 ● (-4)^(n-1)

OpenStudy (solomonzelman):

now, \(a_n=2\cdot(-4)^{n-1}\) \(a_2=2\cdot(-4)^{2-1}=2\cdot (-4)^1=-8\) AND \(a_1=2\cdot(-4)^{1-1}=2\cdot (-4)^0=2 \cdot 1 = 2\). ---------------------------------- YES IT HAS TO BE \(a_n=2\cdot(-4)^{n-1}\)

OpenStudy (solomonzelman):

u r done with this question:)

OpenStudy (anonymous):

thank u!

OpenStudy (solomonzelman):

Anytime!